Predict the major organic product of the following

Predict the major organic product of the following
Answer #1
Concepts and reason

Hydrogen halide addition to the alkene leads to an alkyl halide product through the formation of a carbocation intermediate from the alkene. More stable carbocation leads to the major product. This type of addition is known as Markovnikov’s addition.

Fundamentals

Disered product depends on the stability of the carbocation formed from the alkene in the reaction.

Carbocation stability order

tertiary carbocation> secondary carbocation> primary carbocation> methyl

30>20>10>methyl{{\rm{3}}^{\rm{0}}}{\rm{ > }}{{\rm{2}}^{\rm{0}}}{\rm{ > }}{{\rm{1}}^{\rm{0}}}{\rm{ > }}\,{\rm{methyl}}

Markovnikov’s addition: Negative addendum part in the hydrogen halide is added to the carbon, which is surrounded by fewer hydrogen atoms in the alkene.

Generation of carbocation is as follows

1.Alkene conversion to carbocation by abstraction of a proton from the HI{\rm{HI}}

H&C
ECH₂
CH2 + 8
+ HAL

2.Alkene conversion to carbocation by abstraction of a proton from the HCl{\rm{HCl}}

. من)
+ H
CI
.
CI

3. Alkene conversion to carbocation by abstraction of a proton from the HBr{\rm{HBr}}

+ HLBr
+
Br

Pi-electrons in the given atom abstract the proton from the hydrogen halide to produce carbocation.

Now, the halide ion attacks the carbocation formed in step 1.

1.In this reaction, the halide ion formed is an iodide ion (I)\left( {{{\rm{I}}^{\rm{ - }}}} \right)

CH3 H
H3C
/
“.
—
.
CH,
+
I
-
н
-с
-с
-н

2.In this reaction, the halide ion formed is a chloride ion (Cl)\left( {{\rm{C}}{{\rm{l}}^{\rm{ - }}}} \right)

CI

3.In this reaction, the halide ion formed is a bromide ion (Br)\left( {{\rm{B}}{{\rm{r}}^{\rm{ - }}}} \right)

+
Br

Ans:

The complete reactions are

H
CH2
+
H
-||
HAC-
H
|C
|C
|H
H
+
H
-CT
一一一一
了……
_
#HK)
+ HBr